OpenStudy (anonymous):

If f(x)=x/9sqrt9+x^2 find f '(x) using derivatives

OpenStudy (fibonaccichick666):

just to clarify is this the eq. \[f(x)=\frac{x}{9 \sqrt 9}+x^2\]

OpenStudy (anonymous):

sorry my pc died

OpenStudy (anonymous):

f(x)= x /(9)sqrt(9+x^2))

OpenStudy (fibonaccichick666):

it's cool my internet does it all the time

OpenStudy (fibonaccichick666):

ok so \(f(x)=\frac{x}{9 \sqrt (9+x^2)}\)

OpenStudy (anonymous):

yes

OpenStudy (science0229):

If you simplify the equation, it becomes \[\frac{ 1 }{ 27 }x+x ^{2}\]

OpenStudy (anonymous):

i got (1/9) [(1)(9+x^2)]-(x)92x/sqrt9+x^2)/(9+x^2)^2

OpenStudy (anonymous):

it should say ....-(x)(2x/sqrt..

OpenStudy (psymon):

Just to clarify, we have x[(9√(9+x^2)] ?

OpenStudy (fibonaccichick666):

I wrote it up top Psymon

OpenStudy (anonymous):

x/ [(9√(9+x^2)]

OpenStudy (fibonaccichick666):

so in order to solve this you have two main options, one you can reduce the fraction or two you can apply the quotient rule

OpenStudy (fibonaccichick666):

I am going to assume, the quotient rule is what your teacher is looking for

OpenStudy (fibonaccichick666):

So can you type it for me please?

OpenStudy (anonymous):

??? f'g-fg'/g^2 ?

OpenStudy (psymon):

Yep yep

OpenStudy (anonymous):

(1/9) [(1)(9+x^2)]-(x)92x/sqrt9+x^2)/(9+x^2)^2 <-- i think its right

OpenStudy (fibonaccichick666):

Not quite

OpenStudy (fibonaccichick666):

So in our case, what is g(x)?

OpenStudy (anonymous):

-(x)(2x... I made a typing mistake

OpenStudy (fibonaccichick666):

Still not quite, but what is f(x) as well?

OpenStudy (anonymous):

that it what confuses me since the demoninator says 9sqrt9+x^2 I am assuming that g shall be (9+x^2)

OpenStudy (fibonaccichick666):

that is where the confusion lies

OpenStudy (fibonaccichick666):

The g(x) is the entire bottom piece of the fraction so it is \(g(x)=9 \sqrt(9+x^2)\)

OpenStudy (fibonaccichick666):

So next can you please find g'(x) showing all of your steps?

OpenStudy (anonymous):

ooooh (9)(1/2)(9+2x)? since the 9s dont have an x should I write that they =0?

OpenStudy (fibonaccichick666):

Not quite. Let me show you a similar ex. \[h(x)=8\sqrt(8+x^3)\]\[Let~u=8+x^3~~du=3x^2\]\[Then~we~have: ~~h(x)=8\sqrt u~du\]

OpenStudy (fibonaccichick666):

Whoops forget the du on the last part

OpenStudy (fibonaccichick666):

(also make the last part h(u)) ok so now, can you find the derivative of h(u)?

OpenStudy (anonymous):

it should say 8/sqrt h(u)? is that what you meant?

OpenStudy (anonymous):

you are appliying the chin rule, right?

OpenStudy (fibonaccichick666):

no not h(u) I changed the variable, but yes chain rule.

OpenStudy (fibonaccichick666):

it should read \(h(u)=8\sqrt u\)

OpenStudy (anonymous):

8/sqrtu(u^1/2)

OpenStudy (fibonaccichick666):

ok, not quite

OpenStudy (anonymous):

:(

OpenStudy (fibonaccichick666):

what is the derivative of √u?

OpenStudy (anonymous):

1/2u ^-1/2

OpenStudy (fibonaccichick666):

right

OpenStudy (fibonaccichick666):

so d/dx(8√u)=?

OpenStudy (anonymous):

8(1/2)u^-1/2

OpenStudy (anonymous):

4u^-1/2

OpenStudy (fibonaccichick666):

yes! but one missing piece I asked for the derivative WRT x, so that means ...

OpenStudy (fibonaccichick666):

?